OpenStudy (anonymous):

A ball is thrown straight upward and returns to the thrower’s hand after 4.2 s in the air. A second ball is thrown at an angle of 38◦ with the horizontal. At what speed must the second ball be thrown so that it reaches the same maximum height as the first? The acceleration of gravity is 9.8 m/s^2. Answer in units of m/s

OpenStudy (isaiah.feynman):

@thewinterfawn

OpenStudy (anonymous):

um.. I don't know... that's why I'm asking. ^-^

OpenStudy (anonymous):

I think it might be 0.

OpenStudy (anonymous):

Can you show your work?

OpenStudy (anonymous):

Oh.. okay. thanks (:

OpenStudy (isaiah.feynman):

That's not all!

OpenStudy (anonymous):

It's not? O.o

OpenStudy (anonymous):

How do you find the speed?

OpenStudy (john_es):

Do you still need help?

OpenStudy (anonymous):

Yes please. (:

OpenStudy (john_es):

Well. The first ball cannot be thrown with 0 initial velocity. The ball would fall to the floor in that case.

OpenStudy (john_es):

So, it is needed to find the maximum height of the first ball, but knowing that there should be some initial velocity.

OpenStudy (john_es):

Do you understand that?

OpenStudy (john_es):

Well, to find the initial velocity of the first ball, we use this equation, \[v=v_0-gt\]The ball needs 4.2 s to come back. So the maximum height, the place where the ball has v=0, it reached at 4.2/2=2.1 s. So, \[0=v_0-9.8\cdot2.1\Rightarrow v_0=20.58\ m/s\]

OpenStudy (john_es):

Now, we know the initial velocity of the first ball. So we can find the height that the first ball can reach, \[h=h_0+v_0t-\frac{1}{2}gt^2\]The maximum height is reached at 2.1 s, so, \[h=0+20.58\cdot2.1-4.9\cdot (2.1)^2=21.61\ m\]

OpenStudy (john_es):

Now, we know the maximum height. We can apply that, when the throw is made with an angle, \[v_{0x}=v_0\cos\theta\]\[v_{0y}=v_0\sin\theta\]\[y=y_0+v_{0y}t-\frac{1}{2}gt^2\] But we need to find the time when the ball reach that height, \[v_y=v_{0y}-gt\]

OpenStudy (john_es):

We apply our values, \[21.61=v_{0}\sin(38)t-4.9t^2\] And \[0=v_0\sin(38)-9.8t\Rightarrow t=v_0\sin(38)/9.8=0.0628\cdot v_0\] So, using the first equation, \[21.61=0.0386775v_0^2-0.0193248v_0^2\]

OpenStudy (john_es):

Solving, \[v_0=33.42\ m/s\]

OpenStudy (john_es):

That is the final answer ;).

OpenStudy (john_es):

Do you have the final answer to check it?

OpenStudy (anonymous):

Oops sorry I was doing my calculus homework. Thank you!

OpenStudy (john_es):

No problem, I post the answer in various replies, because the navigator crash eventually :).